# Semi-Autonomous Formalization of the Vlasov-Maxwell-Landau Equilibrium

### 16 Mar 2026 v1

@@ 2.2 Main Result
Theorem 1 (Informal).
@@ 8 Limitations
The theorem – that every smooth positive steady state of the full VML system with Coulomb collisions on the torus must be a global Maxwellian – does not appear in the literature as a stated and proved result. Each proof technique is individually classical: the H-theorem and null-space characterization of the Landau operator are standard [Villani, 2002, cercignani1988]; the polynomial-matching argument that forces constant temperature and Killing’s equation on the bulk velocity appears in cercignani1988 and Desvillettes and Villani [2005]; and the maximum-principle argument for the nonlinear Poisson–Boltzmann equation is textbook. However, prior work assembles these ingredients only for simpler systems. Desvillettes and Villani [2005] prove steady-state uniqueness for the spatially inhomogeneous Boltzmann equation without electromagnetic fields ($E=B=0$). Guo [2003, 2012] study the full VML and Vlasov–Poisson–Landau systems but prove dynamical stability of small perturbations near an assumed Maxwellian, embedding the relevant macroscopic constraints inside linearized energy estimates rather than stating static uniqueness as a standalone result. What is new is the complete static characterization for the electromagnetically coupled system: the observation that the Lorentz force contributes only $\mathcal{O}(|v|^{1})$ terms to the Vlasov equation for a local Maxwellian (because $(v\times B)\cdot v=0$), so the $\mathcal{O}(|v|^{3})$ and $\mathcal{O}(|v|^{2})$ constraints of cercignani1988 and Desvillettes and Villani [2005] carry over unchanged; the Ampère-law integration argument that forces the bulk velocity to vanish; and the closure via Gauss’s law with a neutralizing background. These steps are individually straightforward, but their assembly into a single self-contained proof for the full VML system appears to be new.

### 31 Mar 2026 v2

@@ 2.2 Main Result
Theorem 1 (Informal; see Ilin and Hu [2026] for the complete proof).
@@ 8 Limitations
The theorem – that every smooth positive steady state of the full VML system with Coulomb collisions on the torus must be a global Maxwellian – does not appear in the literature as a stated and proved result. Each proof technique is individually classical: the H-theorem and null-space characterization of the Landau operator are standard [Villani, 2002, Cercignani, 1988]; the polynomial-matching argument that forces constant temperature and Killing’s equation on the bulk velocity appears in Cercignani [1988] and Desvillettes and Villani [2005]; and the maximum-principle argument for the nonlinear Poisson–Boltzmann equation is textbook. However, prior work assembles these ingredients only for simpler systems. Desvillettes and Villani [2005] prove steady-state uniqueness for the spatially inhomogeneous Boltzmann equation without electromagnetic fields ($E=B=0$). Guo [2003, 2012] study the full VML and Vlasov–Poisson–Landau systems but prove dynamical stability of small perturbations near an assumed Maxwellian, embedding the relevant macroscopic constraints inside linearized energy estimates rather than stating static uniqueness as a standalone result. What is new is the complete static characterization for the electromagnetically coupled system: the observation that the Lorentz force contributes only $\mathcal{O}(|v|^{1})$ terms to the Vlasov equation for a local Maxwellian (because $(v\times B)\cdot v=0$), so the $\mathcal{O}(|v|^{3})$ and $\mathcal{O}(|v|^{2})$ constraints of Cercignani [1988] and Desvillettes and Villani [2005] carry over unchanged; the Ampère-law integration argument that forces the bulk velocity to vanish; and the closure via Gauss’s law with a neutralizing background. These steps are individually straightforward, but their assembly into a single self-contained proof for the full VML system appears to be new; the complete proof is given in Ilin and Hu [2026].